HETAV PATEL

Grade 11,

enthalpy means

enthalpy means

Grade:8

4 Answers

vierandra kumar
10 years ago
IT IS THE TOTAL HEAT CONTAING IN THE SYSTEM , IT IS REPRESNTED BY H H=U+PV U=INTENAL ENEGY P= PRESSURE V= VOLUM
Ruchi Shaw
10 years ago
Enthalpy is the total heat of a thermodynamic system.It is equal to the internal energy of the system plus the product of pressure and volume.
Thanks & Regards
Ruchi Shaw
askIITians Faculty
Venkateswarareddy
10 years ago
the sum of internal energy of the system is equal to product of pressure &volume H=E+PV
vierandra kumar
10 years ago
right